# The dual role of constructive agents in public goods games: limited alone, amplifying cooperation with destructive agents

### v1

@@ Abstract
Recent studies have revealed a paradoxical phenomenon in public goods games, wherein destructive agents, harming both cooperators and defectors, can unexpectedly bolster cooperation. Building upon this intriguing premise, our paper introduces a novel concept: constructive agents, which confer additional benefits to both cooperators and defectors. We investigate the impact of these agents on cooperation dynamics within the framework of public goods games. Employing replicator dynamics, we find that unlike destructive agents, the mere presence of constructive agents does not significantly alter the defective equilibrium. However, when the benefits from constructive agents are outweighed by the damage inflicted by destructive agents, the coexistence of both agent types can amplify the ability of destructive agents to sustain cooperation. This occurs through the facilitation of a coexistence pattern involving cooperators, defectors, constructive agents, and destructive agents via cyclic dominance. Conversely, if the benefits from constructive agents surpass the harm caused by destructive agents, the presence of constructive agents nullifies the ability of destructive agents to foster cooperation. Our results highlights the nuanced dual role of constructive agents in shaping cooperation dynamics, emphasizing the necessity of carefully assessing incentive balances in evolutionary game theory.
Evolutionary game theory; Constructive agents; Destructive agents; Public goods game
@@ I Introduction
Cooperation can ensure the maximization of collective interests, but it is also vulnerable to exploitation by “free-riders” who do not bear the costs of cooperation axelrod1981evolution; riolo2001evolution; archetti2012game. To explore the emergence and maintenance of cooperation, numerous studies rooted in the framework of evolutionary game theory have revealed the mechanisms that support the evolution of cooperation in individual interactions weibull1997evolutionary; nowak2011supercooperators; rand2013human; nowak2006five. These include direct reciprocity, established through repeated interactions schmid2021unified; indirect reciprocity involving information transfer leimar2001evolution; and network reciprocity arising from interaction structures nowak2006five. However, in one-shot games where individuals interact only once, without the possibility of memory or behavioral information transfer, these reciprocity mechanisms are absent, making the maintenance of cooperation still a challenge riolo2001evolution. To tackle cooperation challenges in one-shot games, researchers have incorporated social mechanisms that allow individuals to engage in additional decision-making stages after contributing to public goods in the public goods game (PGG), unveiling pathways for establishing cooperation. nowak2011supercooperators; fehr2000cooperation; balliet2011reward; szolnoki2010reward. These include allowing individuals to impose costly punishments fehr2000cooperation; li2018punishment; szolnoki2017second and rewards andreoni2003carrot; chen2015first on others based on contribution outcomes, deciding whether to participate in the game before contributing to public goods (or called “loner” strategy) szabo2002phase; szabo2002evolutionary, or exiting from the interaction early for a fixed payoff shen2021exit; li2024granting. From these game interactions, different behaviors can be observed. The framework of social value orientation analyzes and summarizes these behaviors by exploring how individuals balance their own interests with those of othersbogaert2008social; murphy2011measuring, aiding in understanding cooperation issues in social dilemmas from a perspective of behavioral value orientation.
@@
Recently, attention has been drawn to the impact of sadistic-orientation behavior on the cooperation dynamics, known as destructive behavior (or called “joker” strategy) arenas2011joker; requejo2012stability; khatun2024stability. These “bad guys” not only opt out of contributing to public goods, but also engage in antisocial behaviors that harm the interests of others. Counterintuitively, the presence of these destructive agents can maintain cooperation through a cycle of dominance involving cooperation, defection, and destructive strategies arenas2011joker, and these individuals’ positive effects on cooperation remain robust in both infinite and finite populations requejo2012stability. However, in contrast to this sadistic-orientation behavior, altruistic-orientation behavior is more commonly observed in behavioral experiments bogaert2008social; grund2013natural. For instance, some individuals exhibit a preference for engaging in altruistic rewarding behavior, even at a personal cost choi2013strategic. The complexity of human behavior leads to diversity in actions and strategies in interactions bogaert2008social; murphy2011measuring. Taking this into account, research further explores the evolutionary dynamics of coexisting altruistic and antisocial behaviors gneezy2012conflict. Interestingly, Szolnoki et al. szolnoki2017second demonstrated that antisocial punishment (punishing those who contribute to the public good) can unexpectedly enhance the effectiveness of prosocial punishment (punishing free-riders) in structured populations, challenging conventional perspectives on prosocial/antisocial behaviors and revealing new pathways for establishing cooperation. Importantly, the dynamics created by diverse interaction strategies provide a crucial perspective for a deeper understanding of the emergence of cooperation in real-life scenarios, which lie in the complexity of human behavior santos2012role; perc2008social.
@@
Inspired by this, we introduce a new concept - constructive strategies, incorporating this altruistic behavior into the PGG model involving cooperation, defection, and destructive strategies. This approach enables us to expand current research in solving the cooperation issues by exploring the interconnected diverse behaviors of cooperation dynamics. In contrast to the destructive agents, constructive agents withdraw from game interactions but provide additional costly benefits to participants without concern for their own gains or losses. Utilizing replicator dynamics in an infinite population smith1982evolution, our research reveals that, although constructive agents alone do not alter evolutionary dynamics, their presence alongside destructive agents forms a cycle of dominance involving four strategies that support the emergence of cooperation. Additionally, constructive agents can amplify the role of destructive agents in promoting cooperation. However, when the benefits provided by constructive agents surpass the harm caused by destructive agents, it can result in negative effects that diminish the capacity of destructive agents to promote cooperation. The dual role of constructive individuals emphasizes the balancing impact of positive and negative incentive factors in evolutionary game dynamics, thereby enhancing the understanding of the incentive forces shaping cooperation.
@@ II Model
To gain a deep understanding of how constructors influence the cooperation dynamics, we first examine a three-strategy model that includes cooperation, defection, and construction within the PGG. We then explore a four-strategy PGG model incorporating cooperation, defection, construction, and destruction to study the evolutionary dynamics of cooperation in the presence of both constructive agents and destructive agents.
@@ II.1 Public goods game with constructive agents
In a classic PGG involving $N$ players, a cooperator contributes $c$ to the public pool (simplify without loss of generality, set $c=1$), while a defector contributes nothing. The total amount in the public pool is multiplied by a synergy factor $r$ and then evenly distributed among all participants. Defection is the only Nash equilibrium strategy when the synergy factor for cooperators meets $1<r<N$ archetti2012game. When constructive agents are introduced, these individuals neither contribute to nor benefit from the public pool. Instead, they offer an additional benefit of $d_{2}>0$ to the participants (i.e., cooperators and defectors). Let the number of cooperators among the other individuals in the group be $N_{C}$, and the number of constructive agents be $N_{CA}$, satisfying $0\leq N_{C}+N_{CA}\leq N$, the payoff for a focal individual adopting cooperation ($C$), defection ($D$), and constructive strategies ($CA$) can be respectively expressed as follows:
@@
By utilizing replicator dynamics smith1982evolution, and taking into account the existence of mutations where the probability of an individual mutating to another type is $\mu$ ($\mu\ll 1$), the evolutionary dynamics of strategies in this population can be write as:
@@
where $\bar{P}=xP_{C}+yP_{D}+zP_{CA}$ is the average expected payoffs of the population.
@@ II.2 Public goods game with constructive agents and destructive agents
When considering the four-strategy model, destructive agents (denoted as $DA$), like $CA$, do not participate in public pool allocation and investment, but indiscriminately reduce the payoffs of cooperators and defectors by $d_{1}$ arenas2011joker. Let $N_{DA}$ represent the number of destructive agents, and $S_{2}$ denote the individuals who are neither constructive agents nor destructive agents, i.e., $S_{2}=N-N_{CA}-N_{DA}$. Similar to Eq. (1), the payoffs of focal players adopting $C$, $D$, $CA$, and $DA$ strategies within a group can be expressed as follows:
@@
where $\bar{P}=xP_{C}+yP_{D}+zP_{CA}+wP_{DA}$ is the average expected payoffs of the population.
@@ III Results
Before presenting our research results, we review the dynamics of the destructive agents in the traditional public goods game arenas2011joker; requejo2012stability. In a population of size $N$, assuming that $M$ individuals hold the same strategy. The invasion analysis in ref. arenas2011joker indicates that the destructive agents can lead to three patterns of invasion under the parameter conditions $1<r<r_{\max}=N(M-1)/(M-N)$ and $d_{1}>0$. (i) In the region of $r>1+(N-1)d_{1}$, destructive agents support the emergence of cooperators trough a rock-paper-scissors (RPS) cyclic dominance emerges. (ii) In the region of $1+d_{1}/(M-1)<r<1+(N-1)d_{1}$, a bistability arises between the destructive agents-cooperators. Here, the destructive agents can prevail over the defector, while the defector dominates the cooperator. However, neither the destructive agents nor the cooperator can invade each other. (iii) In the region of $r<1+d_{1}/(M-1)$, the destructive agents dominates in population.
When introducing the constructive agents solely, a similar invasion analysis is conducted under the conditions $0<r<r_{mx}$ and $d_{2}>0$, see Appendix for detailed analysis. The population’s invasion state in the presence of constructive agents is entirely different from that introduced by destructive agents. Figure 1 illustrates that the population consistently favors $D$, since $D$ consistently maintain a positive expected payoff greater than the other two types. But the competitive relationship between $C$ and $CA$ is distinct. Region I: when $r<1-(N-1)d_{2}$ , $C$ can not invade constructive agents, but $CA$ agents can invade $C$, as the expected payoff for the constructive agents is higher than that of cooperators. Region II: When $1-(N-1)d_{2}<r<1-d_{2}/(M-1)$, $C$ and $CA$ can invade each other, leading to a state of antagonism. Region III: when $r>1-d_{2}/(M-1)$, $C$ agents can invade $CA$, since the expected payoff of $C$ is greater than that of $CA$. To validate the theoretical analysis, numerical simulations were conducted in three regions, as illustrated in Figure 2. Consistent with theoretical predictions, across the three mentioned parameter regions, there are variations in the competitive dynamics between $C$ and $CA$, with $D$ consistently holding dominance. These results indicate that when $CA$ exists alone, despite potentially benefiting others, it cannot effectively resolve the social dilemma.
@@
Figure 2: The evolutionary dynamics of a a large, well-mixed population involving strategies $C$, $D$, and $CA$ are derived from the replicator-mutator dynamics. Each vertex in the plot represents a homogeneous population where all individuals hold the same strategy. The three vertices correspond to homogeneous populations adopting the strategies of cooperation, defection, and constructive agents, respectively. Arrows indicate the direction of dynamics, while blue-purple dots indicate the convergence states of populations. Proximity to a vertex indicates a higher proportion of the strategy represented by that vertex. Figures (a), (b), and (c) respectively depict the evolutionary dynamics in region I (setting $r=0.35$ and $d_{2}=0.125$), Region II (setting $r=0.35$ and $d_{2}=1.5$), and Region III (setting $r=3.5$ and $d_{2}=1.5$) as shown in Figure 1. The results presented are generated using a modified version of the DeFinetti package archetti2020definetti.
@@
Figure 3: Depicted are the strategy coexistence in the four-strategy model (involving $C$, $D$, $DA$, and $CA$) as a function of the damage $d_{1}$ caused by destructive agents and the benefit $d_{2}$ provided by constructive agents. The red vertical axis on the far right of each subfigure illustrates the coexistence of strategies in the three-strategy model (involving $C$, $D$, and $DA$) under the same parameters. The square brackets denote the presence of cyclic dominance among strategies, while their absence indicates the state of stable coexistence. The values of $r$ are 3.5, 2, 1, and 0.5 for (a)-(d) respectively, with a group size of $N=5$ and a mutation rate of $\mu=0.005$.
@@
To further explore the impact of constructive agents on cooperation, we investigate the evolutionary dynamics involving the $C$, $D$, $DA$, and $CA$ strategies. The coexistence of strategies obtained through numerical simulation of replicator-mutator equations is depicted in Figure 3. Both the three-strategy model involving $C$, $D$ and $DA$ strategy and the four-strategy model considered here demonstrate that cooperation occurs only when the synergy factor is sufficiently large (i.e., $r\geq 2$), as shown in Figure 3(a) and (b). In the four-strategy model, for a specific $d_{2}$, as the “harm” inflicted by $DA$ on participants ($d_{1}$) increases, the population transitions from a pure $D$ (blue region) phase to a coexistence phase of $D+DA+CA$ (green region). Furthermore, as $d_{1}$ continues to increase, the population shifts from the $D+DA+CA$ phase to the $DA+CA$ (orange region) phase when $r$ is low, as illustrated in Figure 3(c) and (d). Interestingly, a narrow region of four-strategy coexistence (yellow region) emerges between the $D+DA+CA$ phase and the $DA+CA$ phase when $r$ is sufficiently large. In contrast to the three-strategy model (shown on the right red vertical axis in Figure 3) where cooperation is limited to low values of $d_{1}$ arenas2011joker, the introduction of $CA$ broadens the range of $d_{1}$ values conducive to cooperation. Cooperation can persist in regions with large $d_{1}$ as long as $d_{1}>d_{2}$. On the other hand, for a specific $d_{1}$, increasing the benefits provided by $CA$ sustains cooperation within the population. However, as $d_{2}$ surpasses $d_{1}$, cooperation diminishes, leading to a transition from the four-strategy coexistence phase to the $D+DA+CA$ phase. Further increases in $d_{2}$ ultimately lead to the dominance of strategy $D$.
@@
Figure 5: Depicted are the frequency of the $C$ (cooperator, red solid line), $D$ (defector, yellow dashed line), and $DA$ (destructive agents, green dotted line) as a function of time steps. $d_{1}$ is set to 3, 1.65, 1.65, and 4.5 for (a)-(d), respectively. The other parameters are set as follows: $N=5$, $r=3.5$, and $\mu=0.005$. In (a), the frequency of $C$ briefly increases before giving way to $D$. Strategy $DA$ dominates the entire population eventually. In (b), strategies $C$, $D$, and $DA$ exhibit cyclic dominance. The evolutionary process of strategy proportions in (c) and (d) resemble that of Figure (a), with the eventual dominance of strategy $DA$.
Figure 4 illustrates the temporal evolution of four strategies under a synergy factor of $r=3.5$, providing insight into the evolving dynamics within the population. In Figure 4(a), the results indicate a transient increase in the proportions of cooperators and defectors in the early stages of evolution. Subsequently, both $C$ and $D$ decrease while the proportions of $CA$ and $DA$ increase until stability is reached, ultimately resulting in a coexistence state of $CA+DA$. Due to the absence of competition between the $CA$ and $DA$, their proportions in the population are equal at equilibrium. When $d_{2}$ is increased to 1.65, the population exhibits cyclic oscillations over time for these four strategies, as shown in Figure 4(b). Specifically, the proportion of $D$ initially rises, then gives way to $DA$ and $CA$, after which both are gradually replaced by $C$. Then, $C$ is once again supplanted by $D$, forming a cycle. It is this dominating cycle that enables the coexistence of cooperators with the other three strategies. When $d_{1}=d_{2}$, as depicted in Figure 4(c), the dominance of the strategy cycle gives way to the coexistence of $D+DA+CA$. Finally, when the benefits brought by $CA$ outweigh the damage caused by $DA$ (i.e., $d_{2}>d_{1}$), Figure 4(d) shows that defectors eventually dominate the population. In contrast, we examine the temporal evolution dynamics involving three strategies—$C$, $D$, and $DA$. It is noteworthy that when $d_{1}<1$, the cycle dominance can emerge with $C$, $D$, and $DA$, as shown in Figure 5(b). However, when $d_{1}>1$, the results in Figures 5(a), (c), and (d) show that cycle dominance disappears and $DA$ ultimately comes to dominate. Overall, when both $d_{1}$ and $d_{2}$ exceed 1, and satisfy $d_{1}>d_{2}$, the cyclic dominance among strategies can occur (see Figure 4). Yet, with the increasing benefits from $CA$ in the population, the cyclic dominance among the four strategies disappears, indicating that the introduction of constructive factors weakens the ability of destructive factors to maintain cooperation within the population.

### v2

# Constructive agents nullify the ability of destructive agents to foster cooperation in public goods games

@@ Abstract
Existing studies have revealed a paradoxical phenomenon in public goods games, wherein destructive agents, harming both cooperators and defectors, can unexpectedly bolster cooperation. Building upon this intriguing premise, our paper introduces a novel concept: constructive agents, which confer additional benefits to both cooperators and defectors. We investigate the impact of these agents on cooperation dynamics within the framework of public goods games. Employing replicator dynamics, we find that unlike destructive agents, the mere presence of constructive agents does not significantly alter the defective equilibrium. However, when the benefits from constructive agents are outweighed by the damage inflicted by destructive agents, the addition of constructive agents does not affect the ability of destructive agents to sustain cooperation. In this scenario, cooperators can be maintained through a cyclic dominance between cooperators, defectors, and destructive agents, with constructive agents adding complexity but not fundamentally changing the equilibrium. Conversely, if the benefits from constructive agents surpass the harm caused by destructive agents, the presence of constructive agents nullifies the ability of destructive agents to foster cooperation. Our results highlight the nuanced role of constructive agents in cooperation dynamics, emphasizing the necessity of carefully assessing incentive balances when encouraging cooperation.
preprint: AIP/123-QED
Voluntary participation has been shown to be a critical mechanism that promotes cooperation in one-shot and anonymous games. Recent studies have proposed a variant of voluntary participation strategies called destructive agents, who do not participate in the game but harm public goods, which can unexpectedly sustain cooperation. We propose a parallel concept to destructive agents, termed constructive agents, who choose to exit the public goods game but provide additional benefits to public goods. We investigate how these individuals impact the dynamics of cooperation in an infinite and well-mixed population. The results indicate that constructive agents alone cannot alter the equilibrium of defection. However, when they coexist with destructive agents, they nullify the ability of the latter to maintain cooperation when the benefit exceeds the harm to public goods. Our study demonstrates the importance of carefully assessing the balance between positive and negative incentives in alleviating social dilemmas.
@@ I Introduction
Cooperation can ensure the maximization of collective interests, but it is also vulnerable to exploitation by “free-riders" who do not bear the costs of cooperation Axelrod and Hamilton (1981); Riolo, Cohen, and Axelrod (2001); Archetti and Scheuring (2012). To explore the emergence and maintenance of cooperation, numerous studies rooted in the framework of evolutionary game theory have revealed the mechanisms that support the evolution of cooperation in individual interactions Weibull (1997); Nowak and Highfield (2011); Rand and Nowak (2013); Guo et al. (2024); Ren et al. (2024); Nowak (2006). These include direct reciprocity, established through repeated interactions Schmid et al. (2021); indirect reciprocity involving information transfer Leimar and Hammerstein (2001); and network reciprocity arising from interaction structures Nowak (2006). However, in one-shot games where individuals interact only once, without the possibility of memory or behavioral information transfer, these reciprocity mechanisms are absent, making the maintenance of cooperation still a challenge Riolo, Cohen, and Axelrod (2001). To tackle cooperation challenges in one-shot games, researchers have incorporated social mechanisms that allow individuals to engage in additional decision-making stages after contributing to public goods in the public goods game (PGG), unveiling pathways for establishing cooperation Nowak and Highfield (2011); Fehr and Gächter (2000); Balliet, Mulder, and Van Lange (2011); Szolnoki and Perc (2010). These include allowing individuals to impose costly punishments Fehr and Gächter (2000); Li et al. (2018); Szolnoki and Perc (2017) and rewards Andreoni, Harbaugh, and Vesterlund (2003); Chen et al. (2015) on others based on contribution outcomes, deciding whether to participate in the game before contributing to public goods (or called “loner” strategy) Szabó and Hauert (2002a, b), or exiting from the interaction early for a fixed payoff Shen et al. (2021); Li et al. (2024). From these game interactions, different behaviors can be observed. The framework of social value orientation analyzes and summarizes these behaviors by exploring how individuals balance their own interests with those of othersBogaert, Boone, and Declerck (2008); Murphy, Ackermann, and Handgraaf (2011), aiding in understanding cooperation issues in social dilemmas from a perspective of behavioral value orientation.
@@
Recently, attention has been drawn to the impact of sadistic-orientation behavior on the cooperation dynamics, known as destructive behavior (or called “joker” strategy) Arenas et al. (2011); Requejo et al. (2012); Khatun et al. (2024). These “bad guys” not only opt out of contributing to public goods, but also engage in antisocial behaviors that harm the interests of others. Counterintuitively, the presence of these destructive agents can maintain cooperation through a cycle of dominance involving cooperation, defection, and destructive strategies Arenas et al. (2011), and these individuals’ positive effects on cooperation remain robust in both infinite and finite populations Requejo et al. (2012). However, in contrast to this sadistic-orientation behavior, altruistic-orientation behavior is more commonly observed in behavioral experiments Bogaert, Boone, and Declerck (2008); Grund, Waloszek, and Helbing (2013). For instance, some individuals exhibit a preference for engaging in altruistic rewarding behavior, even at a personal cost Choi and Ahn (2013). The complexity of human behavior leads to diversity in actions and strategies in interactions Bogaert, Boone, and Declerck (2008); Murphy, Ackermann, and Handgraaf (2011). Taking this into account, research further explores the evolutionary dynamics of coexisting altruistic and antisocial behaviors Gneezy and Fessler (2012). Interestingly, Szolnoki et al. Szolnoki and Perc (2017) demonstrated that antisocial punishment (punishing those who contribute to the public good) can unexpectedly enhance the effectiveness of prosocial punishment (punishing free-riders) in structured populations, challenging conventional perspectives on prosocial/antisocial behaviors and revealing new pathways for establishing cooperation. Importantly, the dynamics created by diverse interaction strategies provide a crucial perspective for a deeper understanding of the emergence of cooperation in real-life scenarios, which lie in the complexity of human behavior Santos et al. (2012); Perc and Szolnoki (2008).
@@
Inspired by this, we introduce a new concept—constructive strategy, incorporating this altruistic behavior into the PGG model involving cooperation, defection, and destructive strategies. This approach enables us to expand current research in solving the cooperation issues by exploring the interconnected diverse behaviors of cooperation dynamics. In contrast to the destructive agents, constructive agents withdraw from game interactions but provide additional costly benefits to participants without concern for their own gains or losses. Utilizing replicator dynamics in an infinite population Smith (1982), our research reveals that, although constructive agents alone do not alter evolutionary dynamics, when the benefits from constructive agents are outweighed by the damage inflicted by destructive agents, the addition of constructive agents does not affect the ability of destructive agents to sustain cooperation. However, when the benefits provided by constructive agents surpass the harm caused by destructive agents, it can result in negative effects that diminish the capacity of destructive agents to promote cooperation. The role of constructive individuals emphasizes the balancing impact of positive and negative incentive factors in evolutionary game dynamics, thereby enhancing the understanding of the incentive forces shaping cooperation.
@@ II Model
To gain a deep understanding of how constructive agents influence the cooperation dynamics, we first examine a three-strategy model that includes cooperation, defection, and construction within the PGG. We then explore a four-strategy PGG model incorporating cooperation, defection, construction, and destruction to study the evolutionary dynamics of cooperation in the presence of both constructive agents and destructive agents.
@@ II.1 Public goods game with constructive agents
In a classic PGG involving $N$ players, a cooperator contributes $c$ to the public pool (simplify without loss of generality, set $c=1$), while a defector contributes nothing. The total amount in the public pool is multiplied by a synergy factor $r$ and then evenly distributed among all participants. Defection is the only Nash equilibrium strategy when the synergy factor for cooperators meets $1<r<N$ Archetti and Scheuring (2012). When constructive agents are introduced, these individuals neither contribute to nor benefit from the public pool. Instead, they offer an additional benefit of $d_{2}>0$ to the participants (i.e., cooperators and defectors). Let the number of cooperators among the other individuals in the group be $N_{C}$, and the number of constructive agents be $N_{CA}$, satisfying $0\leq N_{C}+N_{CA}\leq{N-1}$, the payoff for a focal individual adopting cooperation ($C$), defection ($D$), and constructive agents strategies ($CA$) can be respectively expressed as follows:
@@
By utilizing replicator dynamics Smith (1982), and taking into account the existence of mutations where the probability of an individual mutating to another type is $\mu$ ($\mu\ll 1$), the evolutionary dynamics of strategies in this population can be write as:
@@
where $\bar{P}=xP_{C}+yP_{D}+zP_{CA}$ is the expected payoff of the population.
@@ II.2 Public goods game with constructive agents and destructive agents
When considering the four-strategy model, destructive agents (denoted as $DA$), like $CA$, do not participate in public pool allocation and investment, but indiscriminately reduce the payoffs of cooperators and defectors by $d_{1}$ Arenas et al. (2011). Let $N_{DA}$ represent the number of destructive agents, and $S_{2}$ denote the individuals who are neither constructive agents nor destructive agents, i.e., $S_{2}=N-N_{CA}-N_{DA}$. Similar to Eq. (1), the payoffs of focal players adopting $C$, $D$, $CA$, and $DA$ strategies within a group can be expressed as follows:
@@
where $\bar{P}=xP_{C}+yP_{D}+zP_{CA}+wP_{DA}$ is the average expected payoff of the population.
@@ III Results
Before presenting our research results, we review the dynamics of the destructive agents in the traditional public goods game Arenas et al. (2011); Requejo et al. (2012). In a population of size $M$, it is assumed that all individuals hold the same strategy. The invasion analysis in ref. Arenas et al. (2011) indicates that the destructive agents can lead to three patterns of invasion under the parameter conditions $1<r<r_{\max}=N(M-1)/(M-N)$ and $d_{1}>0$. (i) In the region of $r>1+(N-1)d_{1}$, destructive agents support the emergence of cooperators trough a rock-paper-scissors (RPS) cyclic dominance emerges. (ii) In the region of $1+d_{1}/(M-1)<r<1+(N-1)d_{1}$, a bistability arises between the destructive agents-cooperators. Here, the destructive agents can prevail over the defector, while the defector dominates the cooperator. However, neither the destructive agents nor the cooperator can invade each other. (iii) In the region of $r<1+d_{1}/(M-1)$, the destructive agents dominates in population.
When introducing the constructive agents solely, a similar invasion analysis is conducted under the conditions $0<r<{r_{max}=N(M-1)/(M-N)}$ and $d_{2}>0$, see Appendix A for detailed analysis. The population’s invasion state in the presence of constructive agents is entirely different from that introduced by destructive agents. Figure 1 illustrates that the population consistently favors $D$, since $D$ consistently maintain a positive expected payoff greater than the other two types. But the competitive relationship between $C$ and $CA$ is distinct. Region I: when $r<1-(N-1)d_{2}$ , $C$ can not invade constructive agents, but $CA$ agents can invade $C$, as the expected payoff for the constructive agents is higher than that of cooperators. Region II: When $1-(N-1)d_{2}<r<1-d_{2}/(M-1)$, $C$ and $CA$ can invade each other, leading to a state of antagonism. Region III: when $r>1-d_{2}/(M-1)$, $C$ agents can invade $CA$, since the expected payoff of $C$ is greater than that of $CA$. To validate the theoretical analysis, numerical simulations were conducted in three regions, as illustrated in Figure 2. Consistent with theoretical predictions, across the three mentioned parameter regions, there are variations in the competitive dynamics between $C$ and $CA$, with $D$ consistently holding dominance. These results indicate that when $CA$ exists alone, despite potentially benefiting others, it cannot effectively resolve the social dilemma.
@@
Figure 2: The evolutionary dynamics of a a large, well-mixed population involving strategies $C$, $D$, and $CA$ are derived from the replicator-mutator dynamics. Each vertex in the plot represents a homogeneous population where all individuals hold the same strategy. The three vertices correspond to homogeneous populations adopting the strategies of cooperation, defection, and constructive agents, respectively. Arrows indicate the direction of dynamics, while blue-purple dots indicate the convergence states of populations. Proximity to a vertex indicates a higher proportion of the strategy represented by that vertex. Figures (a), (b), and (c) respectively depict the evolutionary dynamics in region I (setting $r=0.35$ and $d_{2}=0.125$), Region II (setting $r=0.35$ and $d_{2}=1.5$), and Region III (setting $r=3.5$ and $d_{2}=1.5$) as shown in Figure 1. The other parameters are set as follows: $N=5$ and $\mu=0.005$. The results presented are generated using a modified version of the DeFinetti package Archetti (2020).
@@
Figure 3: Depicted are the strategy coexistence in the four-strategy model (involving $C$, $D$, $DA$, and $CA$) as a function of the damage $d_{1}$ caused by destructive agents and the benefit $d_{2}$ provided by constructive agents. The red vertical axis on the far right of each subfigure illustrates the coexistence of strategies in the three-strategy model (involving $C$, $D$, and $DA$) under the same parameters. The square brackets denote the presence of cyclic dominance among strategies, while their absence indicates the state of stable coexistence. Due to mutations, a given strategy cannot completely disappear or dominate; therefore, a threshold of greater than 0.05 is used to determine its presence. The values of $r$ are 3.5, 2, 1, and 0.5 for (a)-(d) respectively, with a group size of $N=5$ and a mutation rate of $\mu=0.005$.
@@
To further explore the impact of constructive agents on cooperation, we investigate the evolutionary dynamics involving the $C$, $D$, $DA$, and $CA$ strategies. The coexistence of strategies obtained through numerical simulation of replicator-mutator equations is depicted in Figure 3. Both the three-strategy model involving $C$, $D$ and $DA$ strategy and the four-strategy model considered here demonstrate that cooperation occurs only when the synergy factor is sufficiently large (i.e., $r\geq 2$), as shown in Figure 3(a) and (b). In the four-strategy model, for a specific $d_{2}$, as the “harm" inflicted by $DA$ on participants ($d_{1}$) increases, the population transitions from a pure $D$ (blue region) phase to a coexistence phase of $D+DA+CA$ (green region). Furthermore, as $d_{1}$ continues to increase, the population shifts from the $D+DA+CA$ phase to the $DA+CA$ (orange region) phase when $r$ is low, as illustrated in Figure 3(c) and (d). Interestingly, a narrow region of four-strategy coexistence (yellow region) emerges between the $D+DA+CA$ phase and the $DA+CA$ phase when $r$ is sufficiently large. In contrast to the three-strategy model (shown on the right red vertical axis in Figure 3) where cooperation is limited to low values of $d_{1}$ Arenas et al. (2011), the introduction of $CA$ enables cooperation to emerge under higher values of $d_{1}$. However, the premise for cooperation to be sustained is that $d_{1}$ must be greater than $d_{2}$. Once $d_{2}$ exceeds $d_{1}$, cooperation diminishes, leading to a transition from the four-strategy coexistence phase to the $D+DA+CA$ phase. Further increases in $d_{2}$ ultimately lead to the dominance of strategy $D$.
@@
Figure 5: Depicted are the frequency of the $C$ (cooperator, red solid line), $D$ (defector, yellow dashed line), and $DA$ (destructive agents, green dotted line) as a function of time steps. $d_{1}$ is set to 3, 0.65, 1.65, and 4.5 for (a)-(d), respectively. The other parameters are set as follows: $N=5$, $r=3.5$, and $\mu=0.005$. In (a), the frequency of $C$ briefly increases before giving way to $D$. Strategy $DA$ dominates the entire population eventually. In (b), strategies $C$, $D$, and $DA$ exhibit cyclic dominance. The evolutionary process of strategy proportions in (c) and (d) resemble that of Figure (a), with the eventual dominance of strategy $DA$.
Figure 4 illustrates the temporal evolution of four strategies under a synergy factor of $r=3.5$, providing insight into the evolving dynamics within the population. In Figure 4(a), the results indicate a transient increase in the proportions of cooperators and defectors in the early stages of evolution. Subsequently, both $C$ and $D$ decrease while the proportions of $CA$ and $DA$ increase until stability is reached, ultimately resulting in a coexistence state of $CA+DA$. Due to the absence of competition between the $CA$ and $DA$, their proportions in the population are equal at equilibrium. When $d_{2}$ is increased to 1.65, the population exhibits cyclic oscillations over time for these four strategies, as shown in Figure 4(b). Specifically, the proportion of $D$ initially rises, then gives way to $DA$ and $CA$, after which both are gradually replaced by $C$. Then, $C$ is once again supplanted by $D$, forming a cycle. It is this dominating cycle that enables the coexistence of cooperators with the other three strategies. When $d_{1}=d_{2}$, as depicted in Figure 4(c), the dominance of the strategy cycle gives way to the coexistence of $D+DA+CA$. Finally, when the benefits brought by $CA$ outweigh the damage caused by $DA$ (i.e., $d_{2}>d_{1}$), Figure 4(d) shows that defectors eventually dominate the population. In contrast, we examine the temporal evolution dynamics involving three strategies—$C$, $D$, and $DA$. It is noteworthy that when $d_{1}<1$, the cycle dominance can emerge with $C$, $D$, and $DA$, as shown in Figure 5(b). However, when $d_{1}>1$, the results in Figures 5(a), (c), and (d) show that cycle dominance disappears and $DA$ ultimately comes to dominate. Overall, when both $d_{1}$ and $d_{2}$ are larger, the cyclic dominance among strategies can occur (see Figure 4), but it must satisfy $d_{1}>d_{2}$. However, as the increasing benefits from $CA$ in the population, the cyclic dominance among strategies induced by destructive agents disappears when $d_{2}\geq d_{1}$, indicating that the introduction of constructive agents weakens the ability of destructive agents to maintain cooperation within the population. The results remain consistent even when constructive and destructive agents are integrated (see Appendix B).